# Hedging Hydrogen: Planning and Contracting Under Uncertainty for a Green Hydrogen Producer

### v2

@@ 4.3.8 Green Hydrogen Subsidies
Green hydrogen subsidies are assumed to be only obtainable with the use of hour-by-hour temporally correlated renewable power (in the context of this study, this means power originating from renewable PPA’s). The subsidy-eligible quantity of hydrogen produced at hour in scenario , is defined with the constraints:
@@ 5 Results and Discussion
The results are summarised in Figure 5, and the out-of-sample test results for each solution in Table 5. The computed Policy Comparison Metrics are shown in Table 6. Additional data on equipment sizes and hedging decisions are provided in the Supplementary Material.
LCOH (€/kg) | ||||
Solution | Mean |
6.70 | 10.69 | |||
6.51 | 7.25 | |||
6.69 | 10.42 | |||
6.57 | 7.06 | |||
6.86 | 6.99 | |||
4.38 | 5.14 | |||
4.23 | 4.70 | |||
7.23 | 9.24 | |||
6.71 | 7.80 | |||
5.49 | 7.50 | |||
4.53 | 5.22 |
@@ 5.1 Fixed Demand HPA
As was the case in [56], the low VSS indicates little to no value is obtained using a stochastic model for the Fixed Demand HPA.
@@
The Value of the Resale-Enabled Solution (VRES) is positive on mean performance (4.3%) and slightly negative on worst-case performance (-1.1%). This is to be expected, because while both policies and are equally risk-averse, the no-resale formulation will systematically underestimate hedging options with good re-sale value, and so it finds less optimal solutions on average.
@@ 5.4 Uncertain Demand HPA – Subsidised
We also see that the Cost of Demand Uncertainty is increased when green subsidies are available: the cost of an Uncertain Demand HPA – Subsidised is a 6.6% increase on average with respect to a Fixed Demand HPA – Subsidised.
@@ 6 Conclusion
Demand uncertainty in HPA contracts was also seen to necessitate increased equipment sizing (of the electrolyser and hydrogen storage). This increased equipment sizing may be a serious obstacle for some project developers facing difficulties in obtaining financing for the initial investment costs, or in obtaining sufficient hydrogen storage on-site. As such, in the negotiation phase of a HPA, both consumer and producer should consider what forms and levels of demand uncertainty are acceptable to both parties, with the knowledge that less strict contracts may necessitate increased sale prices to fund increased equipment CAPEX costs.
@@ 7 Acknowledgements
This study was carried out as part of a collaboration between Mines Paris (PSL University) and Verso Energy. The authors thank the Verso Energy Offers and Partnerships team for providing consultative support, especially Arthur Auxenfants, Jean-Baptiste Martin, and Julien Guiet. The problem is implemented using proprietary modelling software Versys, Verso Energy. Many thanks also to the reviewers of the first submission of this article who provided many detailed and useful recommendations.

### v3

@@ 4.3.8 Green Hydrogen Subsidies
Green hydrogen subsidies are assumed to be only attainable with the use of hour-by-hour temporally correlated renewable power (in the context of this study, this means power originating from renewable PPA’s). The subsidy-eligible quantity of hydrogen produced at hour in scenario , is defined with the constraints:
These constraints effectively perform a max operation between the total PPA power available in a given hour, and the electrolyser production in the same hour. Thus making sure that no more hydrogen can be counted than is actually produced for that hour, and that the amount of green hydrogen counted in that hour cannot exceed the (conversion efficiency-adjusted) renewable power received for that hour from the PPA’s. Additional power from PPA’s which is instead sold to the market in a given hour thus forgoes its ‘green’ credentials.
@@ 5 Results and Discussion
The results are summarised in Figure 6, and the out-of-sample test results for each solution in Table 6. The computed Policy Comparison Metrics are shown in Table 5 and displayed in Figures 7 and 8. Additional data on equipment sizes and hedging decisions are provided in the Supplementary Material.
@@ 5.1 Fixed Demand HPA
As was the case for Palmer2023Risk-conscious., the low VSS indicates little to no value is obtained using a stochastic model for the Fixed Demand HPA.
@@
The Value of the Resale-Enabled Solution (VRES) is positive on mean performance (4.3%) and slightly negative on worst-case performance (-1.1%). This is to be expected, because while both policies and are equally risk-averse, the no-resale formulation will systematically underestimate hedging options with good re-sale value, and so it finds less optimal solutions on average. Note that we are only able to obtain a VRES for the risk-averse case, as for the risk-neutral case energy resale cannot be enabled without rendering the results speculative and nonsensical, as discussed in Section 1.3.
@@ 5.4 Uncertain Demand HPA – Subsidised
We also see that the Cost of Demand Uncertainty is increased when green subsidies are available: the cost of an Uncertain Demand HPA – Subsidised is a 6.6% increase on average with respect to a Fixed Demand HPA – Subsidised (see Figure 8).
@@ 6 Conclusion and Policy Recommendations
Demand uncertainty in HPA contracts was also seen to necessitate increased equipment sizing (of the electrolyser and hydrogen storage). This increased equipment sizing may be a serious obstacle for some project developers facing difficulties in obtaining financing for the initial investment costs, or in obtaining sufficient hydrogen storage on-site. As such, in the negotiation phase of an HPA, both consumer and producer should consider what forms and levels of demand uncertainty are acceptable to both parties, with the knowledge that less strict contracts may necessitate increased sale prices to fund increased equipment CAPEX costs.
@@ 7 Acknowledgements
This study was carried out as part of a collaboration between Mines Paris (PSL University) and Verso Energy. The authors thank the Verso Energy Offers and Partnerships team for providing consultative support, especially Arthur Auxenfants, Jean-Baptiste Martin, and Julien Guiet. The problem is implemented using proprietary modelling software Versys, Verso Energy. Many thanks also to the reviewers of this article who provided many detailed and useful recommendations.
8 Data Availability
All scenarios used for in-sample and out-of-sample testing are available at https://github.com/owen-verso/Hedging-Hydrogen-Data-Package.